OpenStudy (anonymous):

Which of the following is an equation of the line that passes through (-3,5) with a slope of -5 A. y + 5 = -5(x - 3) B. y - 5 = -5(x + 3) C. y - 5 = -3(x - 5) D. y + 5 = -3(x + 5

OpenStudy (anonymous):

I will give a medal!

OpenStudy (imtiaz7):

if you not give me then i'll also help you bree

OpenStudy (anonymous):

can you help me on this?

OpenStudy (imtiaz7):

yes i'm trying to help you

OpenStudy (imtiaz7):

ok it can be solved by point slope form, that is: y-y1=m(x-x1) so here x1=-3 and y1=5 and m=-5 just put the values and solve it

OpenStudy (imtiaz7):

it means "B" is correct one! @bree_ramos i hope you understand :)

OpenStudy (anonymous):

thank you so much!!

OpenStudy (imtiaz7):

It's good that you understand and welcome @bree_ramos